# Soban Nasir Lone

### v1

@@ 3 Proposed $\alpha$-VI DeepONet
We propose using Bayesian neural networks in both the trunk and branch networks (shown in Fig. 2) instead of their deterministic counterparts, similar to VB-DeepONet. However, a fully factorised normal distribution is hardly ideal as a prior for BNNs. Constructing alternative prior beliefs that accurately reflect our judgments could be computationally prohibitive too. Nonetheless, by employing an alternative divergence metric ($D$), more robust posterior beliefs can be produced with an imperfect prior. Therefore, in the proposed approach, we seek to use a divergence metric, different from KLD, that results in posteriors more robust to poorly specified priors and provides reliable uncertainty quantification.
In the deterministic DeepONet, the parameter vector $\bm{\theta}$ of the deep neural network (from Eq. 4) is a point estimate. However, as noted in [32], multiple different parameter settings can perform equally well, implying that the set of values that each parameter can take may be captured by a distribution over those plausible values. A Bayesian paradigm allows us to place distributions on the model parameters, representing uncertainty about their exact values. Using Bayes’ rule, we update these parameter distributions with the data ($\mathcal{D}$ from Eq. 3) as follows:
@@
$p(\mathcal{D}^{*}|\mathcal{D},\bm{{\theta}})=\int p(\mathcal{D}^{*}|\bm{{\theta}})p(\bm{{\theta}}|\mathcal{D})\text{d}\bm{{\theta}}.$ (6)
@@ 3.1 Fully Factorised Prior
where $L$ denotes the dimensionality of the model parameter vector $\bm{{\theta}}$. It is important to acknowledge that this fully factorised prior is a “misspecified prior” as it does not capture the interdependencies that exist between the individual components $\theta_{l}$ of the parameter vector $\bm{{\theta}}$. To alleviate the negative influence of this “misspecified prior,” we use a different divergence metric in Section 3.4.
@@ 3.3 Posterior Approximation via Variational Inference
Within a fully Bayesian framework, our goal is to determine the posterior distribution of the model parameters given the observed data. This posterior distribution, denoted as $p(\bm{{\theta}}|\mathcal{D})$, represents the probability of various parameter configurations ($\bm{{\theta}}$) after observing the data $\mathcal{D}$. Unfortunately, due to the intractable normalising constant as explained in Section 3, directly computing the posterior distribution analytically is not feasible.
Generally, there are two approaches to solve this class of problems - sampling-based approaches, e.g., Markov chain Monte Carlo and approximation-based optimisation approaches, e.g., variational inference, expectation propagation (EP). Sampling-based approaches, exemplified by MCMC-techniques, are often regarded as the gold standard for Bayesian inference [33]. They involve constructing a Markov chain that gradually explores the parameter space, eventually converging to a distribution proportional to the posterior. However, for problems with high-dimensional parameter spaces (often exceeding 20,000 parameters in our case), these approaches can become computationally impractical [34]. We note here that the proposed approach puts no limit on the complexity of the model, that is the number of parameters. Hence, we will focus on the approximation-based approaches, where a simpler variational distribution is introduced to approximate the posterior, and the key lies in optimising the parameters of the variational distribution to minimise the Kullback-Liebler divergence between itself and the intractable posterior. This optimisation process allows us to efficiently train complex models despite the high dimensionality of the parameter space.
@@ 3.3.1 Variational Inference with KLD
VI [35] introduces a surrogate posterior family of distributions, denoted by $Q$, that is tractable to sample from. The goal is to find a member of $Q$ distribution that closely resembles the true posterior in terms of the shape. VI accomplishes this by minimising the KLD between the approximate posterior $q(\bm{{\theta}};\bm{{\eta}})\in Q$ and the true posterior $p(\bm{{\theta}}|\mathcal{D})$. The KLD, denoted as $D_{KL}[q\mid\mid p]$, quantifies the difference between the $q$ distribution and the reference $p$ distribution weighted by the $q$ distribution, and is mathematically defined as:
@@
where the expectation in the ELBO (in Eq. 13) is taken with respect to the approximate posterior $q(\bm{{\theta}};\bm{{\eta}})$. Note that $p(\mathcal{D})$ is the model evidence, which is a constant term for a given dataset and also constant with respect to $q(\bm{{\theta}};\bm{{\eta}})$ and can therefore be ignored during the minimisation process Eq. 12. From Eq. 13, we see that maximising the ELBO is equivalent to minimising the KLD (as KLD is non-negative). So, VI aims to maximise the ELBO which turns out to be the minimisation of the variational free energy $\mathcal{L}(q)$ (see Eq. 14. This translates to finding an approximate posterior $q(\bm{{\theta}};\bm{{\eta}})$ that balances two key factors (common in Bayesian inference):
@@
The expressions of these terms highlight the impact of our prior modelling choices on the optimisation problem. A poorly specified prior – independent standard Gaussian distribution – over parameters, as is the case with BNNs, can degrade the quality of the posterior approximation. However, it remains the most expedient choice for BNNs from the perspective of computational expense. As suggested by Knoblauch et al. [26], we can opt for a difference divergence metric instead of KLD to alleviate the negative influence of misspecified prior. For example, KLD is known to exhibit a mode-seeking behaviour [25, 36]. By selecting a different divergence metric one can improve the quality of the posterior distribution, both in terms of fitting the data and providing better uncertainty estimates [25, 26, 37].
In this context, we examine the effect of using the flexible Rényi’s-$\alpha$ divergence on the posterior predictive performance. We modify the variational free energy objective by replacing the KLD with Rényi’s-$\alpha$ divergence:
$\mathcal{L}(q)=-\mathbb{E}_{q(\bm{{\theta}};\bm{{\eta}})}[\log p(\mathcal{D}|\bm{{\theta}})]+D_{AR}^{(\alpha)}[q(\bm{{\theta}};\bm{\eta})||p(\bm{{\theta}})].$ (15)
To compute the expected value of the log-likelihood term, we use Monte-Carlo integration by drawing samples from the variational distribution $q$ given the parameters $\bm{{\eta}}$. The likelihood function is computed using Section 3.2 and we approximate the first term in Eq. 15, as follows:
@@
[…]
where $N_{q}$ is the number of Monte Carlo samples, and $\bm{{\theta}}^{(c)}$ is the $c^{\text{th}}$ sample drawn from $q(\bm{{\theta}};\bm{\eta})$. In practice, we found that using 25 Monte Carlo samples ($N_{q}=25$) is sufficient for accurate approximation, aligning with findings in [38] which suggest that even a single sample can yield satisfactory results. In the next section, we detail Rényi’s $\alpha$-divergence as an alternative to KLD, which represents the second term of our objective function.
3.4 Rényi’s-$\alpha$ divergence
Rényi’s $\alpha$-divergence, introduced by Rényi [29], offers a robust alternative to KLD. It is characterised by a hyperparameter $\alpha$, which controls the degree of robustness, and it converges to KLD as a limiting case when $\alpha$ approaches 1. We denote this divergence by $D_{AR}^{(\alpha)}$ and follow the parameterisation provided in [39]:
$D_{AR}^{(\alpha)}[q(\theta)||p(\theta)]=\frac{1}{\alpha(\alpha-1)}\log\left(\mathbb{E}_{q(\theta)}\left[\left(\frac{p(\theta)}{q(\theta)}\right)^{1-\alpha}\right]\right).$ (17)
@@ 3.4 Rényi’s-$\alpha$ divergence
In the context of our problem, $D_{AR}^{(\alpha)}$ represents the second term of the objective function in Eq. 15, which is approximated as follows:
$D_{AR}^{(\alpha)}[q(\bm{{\theta}};\bm{\eta})||p(\bm{{\theta}})]\approx\frac{1}{\alpha(\alpha-1)}\log\left(\frac{1}{N_{q}}\sum_{c=1}^{N_{q}}\left[\left(\frac{p(\bm{{\theta}}^{(c)}))}{q(\bm{{\theta}}^{(c)};\bm{\eta})}\right)^{1-\alpha}\right]\right),$ (19)
@@
In Fig. 3, we plot divergence values $D_{AR}^{(\alpha)}$ between $q(\theta)$ and $p(\theta)$, assuming both are normal distributions with $p(\theta)$ as a standard normal distribution. The divergence value decreases as $\alpha$ increases, with KLD as the case when $\alpha=1$. Note that the objective function in Eq. 15 contains a data-loss term and a complexity term. The behaviour of the optimisation problem differs significantly depending on whether $\alpha$ is greater than or less than 1. For $\alpha>1$, $D_{AR}^{(\alpha)}$ exhibits a mode-seeking behaviour, causing the posteriors to become more concentrated while being penalised less for deviating from the prior distribution. In this regime, the objective function places more emphasis on the data-loss term, and the complexity term – which measures the divergence between the prior and the variational posterior – receives less weight as $\alpha$ increases. In contrast, for $\alpha<1$, $D_{AR}^{(\alpha)}$ encourages a mass-covering behaviour, akin to expectation propagation. As $\alpha$ decreases below 1, the values of $D_{AR}^{(\alpha)}$ increase, resulting in more weight being assigned to the complexity term and less to the data-loss term. It is important to note, however, that a larger regularisation term does not necessarily imply that a misspecified prior will dominate the inference, as demonstrated in the experiments by [26].
@@
Given that the true posterior is unknown, it is uncertain whether the mode-seeking or mass-covering behaviour is preferable, making the choice problem-dependent. For scenarios where the true posterior is multimodal, a divergence that produces wider parameter variances (i.e., mass-covering behavior) may be beneficial for achieving more robust predictive estimates. In particular, we aim to set $D_{AR}^{(\alpha)}$ such that the posterior is robust to priors that strongly contradict observed data, and provides reliable uncertainty quantification. We rely on validation sets and adjust $\alpha$ as a hyperparameter to minimise metrics such as the normalised mean squared error or negative log-likelihood.

### v2

@@ 3 Proposed $\alpha$-VI DeepONet
We propose using Bayesian neural networks in both the trunk and branch networks (shown in Fig.˜2) instead of their deterministic counterparts, similar to VB-DeepONet. However, a fully factorised normal distribution is hardly ideal as a prior for BNNs. Constructing alternative prior beliefs that accurately reflect our judgements could be computationally prohibitive too. Nonetheless, by employing an alternative divergence metric ($D$), more robust posterior beliefs can be produced with an imperfect prior. Therefore, in the proposed approach, we seek to use a divergence metric, different from KLD, that results in posteriors more robust to poorly specified priors and provides reliable uncertainty quantification. Aside from this change in divergence, our approach adheres to the same distributional assumptions as VB-DeepONet [garg2023vb], as outlined in the following sections.
In the deterministic DeepONet, the parameter vector $\bm{\theta}$ of the deep neural network (from Eq.˜4) is a point estimate. However, as noted in [choromanska2015loss], multiple different parameter settings can perform equally well, implying that the set of values that each parameter can take may be captured by a distribution over those plausible values. A Bayesian paradigm allows us to place distributions on the model parameters, representing uncertainty about their exact values. Using Bayes’ rule, we update these parameter distributions with the data ($\mathcal{D}$ from Eq.˜3) as follows:
@@
$p(\mathcal{D}^{*}|\mathcal{D})=\int p(\mathcal{D}^{*}|\bm{{\theta}})p(\bm{{\theta}}|\mathcal{D})\text{d}\bm{{\theta}}.$ (6)
@@ 3.1 Fully factorised prior
where $L$ denotes the dimensionality of the model parameter vector $\bm{{\theta}}$. While fully factorised Gaussian priors remain the most expedient choice in practice, they often fail to capture dependencies between parameters and can therefore be regarded as misspecified in many scenarios. This concern has been empirically validated in the context of Bayesian neural networks, where such priors have been linked to the cold posterior effect [wenzel2020good, fortuin2021bayesian]. At the same time, Pearce et al. [pearce2020expressive] demonstrated that more expressive priors can indeed be constructed even when starting from a factorised Gaussian assumption, for example by composing kernels in the induced function space. These complementary perspectives reinforce the importance of either designing richer priors or adopting inference procedures that are robust to prior misspecification. To alleviate the negative influence of this prior, we use a different divergence metric in Section˜3.4.
@@ 3.3 Posterior approximation via variational inference
Within a fully Bayesian framework, our goal is to determine the posterior distribution of the model parameters given the observed data. This posterior distribution, denoted as $p(\bm{{\theta}}|\mathcal{D})$, represents the probability of various parameter configurations ($\bm{{\theta}}$) after observing the data $\mathcal{D}$. Unfortunately, due to the intractable normalising constant as explained in Section˜3, directly computing the posterior distribution analytically is not feasible.
Generally, there are two approaches to solve this class of problems - sampling-based approaches, e.g., Markov chain Monte Carlo and approximation-based optimisation approaches, e.g., variational inference, expectation propagation (EP). Sampling-based approaches, exemplified by MCMC techniques, are often regarded as the gold standard for Bayesian inference [gilks1995markov]. They involve constructing a Markov chain that gradually explores the parameter space, eventually converging to a distribution proportional to the posterior. Among these, Hamiltonian Monte Carlo (HMC) stands out for its efficiency in exploring high-dimensional spaces, using gradient information to propose distant transitions that maintain high acceptance rates. Despite these advantages, HMC is often prohibitively expensive in practice, requiring repeated full-batch gradient evaluations and long burn-in periods. Recent work, such as Izmailov et al. [izmailov2021bayesian], has shown that scalable HMC is possible using massive computational resources (e.g., hundreds of TPUs), but such setups remain impractical for most researchers and applications. Making sampling-based inference tractable and efficient for deep models remains an open and active area of research.
In our case, the problem setup involves high-dimensional parameter spaces, often exceeding 20,000 parameters. For such settings, sampling-based approaches become computationally infeasible. This motivates the use of approximation-based methods, particularly variational inference, which offer a more tractable alternative. VI introduces a simpler variational distribution to approximate the true posterior and reframes inference as an optimisation problem by minimising the Kullback-Leibler divergence between the variational distribution and the true posterior. Unlike MCMC, VI scales naturally with the dimensionality of the parameter space and imposes no explicit constraints on its complexity, allowing efficient training of large-scale Bayesian neural networks.
@@ 3.3.1 Variational inference with KLD
VI [blei2017variational] introduces a surrogate posterior family of distributions, denoted by $Q$, that is tractable to sample from. The goal is to find a member of $Q$ distribution that closely resembles the true posterior in terms of the shape. VI accomplishes this by minimising the KLD between the approximate posterior $q(\bm{{\theta}};\bm{{\eta}})\in Q$ and the true posterior $p(\bm{{\theta}}|\mathcal{D})$. The KLD, denoted as $D_{KL}[q\mid\mid p]$, quantifies the difference between the $q$ distribution and the reference $p$ distribution weighted by the $q$ distribution, and is mathematically defined as:
@@
where the expectation in the ELBO (in Eq.˜13) is taken with respect to the approximate posterior $q(\bm{{\theta}};\bm{{\eta}})$. Note that $p(\mathcal{D})$ is the model evidence, which is a constant term for a given dataset and also constant with respect to $q(\bm{{\theta}};\bm{{\eta}})$ and can therefore be ignored during the minimisation process Eq.˜12. From Eq.˜13, we see that maximising the ELBO is equivalent to minimising the KLD (as KLD is non-negative). So, VI aims to maximise the ELBO which turns out to be the minimisation of the variational free energy $\mathcal{L}(q)$ (see Eq.˜14). This translates to finding an approximate posterior $q(\bm{{\theta}};\bm{{\eta}})$ that balances two key factors (common in Bayesian inference):
@@
The expressions of these terms highlight the impact of our prior modelling choices on the optimisation problem. A poorly specified prior – independent standard Gaussian distribution – over parameters, as is the case with BNNs, can degrade the quality of the posterior approximation. However, it remains the most expedient choice for BNNs from the perspective of computational expense. As suggested by Knoblauch et al. [knoblauch2019generalized], we can opt for a difference divergence metric instead of KLD to alleviate the negative influence of misspecified prior. For example, KLD is known to exhibit a mode-seeking behaviour [li2016renyi, hernandez2015probabilistic]. By selecting a different divergence metric one can improve the quality of the posterior distribution, both in terms of fitting the data and providing better uncertainty estimates [li2016renyi, knoblauch2019generalized, knoblauch2018doubly].
In this context, we examine the effect of using the flexible Rényi’s $\alpha$-divergence on the posterior predictive performance. We modify the objective function by replacing the KLD with Rényi’s $\alpha$-divergence. This objective function, $\mathcal{L}(q)$, represents the variational free energy to be minimised and is a direct application of the Generalized Variational Inference framework ([knoblauch2019generalized], Eq. 10):
$\mathcal{L}(q)=-\mathbb{E}_{q(\bm{{\theta}};\bm{{\eta}})}[\log p(\mathcal{D}|\bm{{\theta}})]+D_{AR}^{(\alpha)}[q(\bm{{\theta}};\bm{\eta})\mid\mid p(\bm{{\theta}})],$ (15)
where $D_{AR}^{(\alpha)}$ is the Rényi $\alpha$-divergence defined in the following section. This replaces $D_{KL}$ from standard VI while keeping $p(\bm{{\theta}})$ and $q(\bm{{\theta}};\bm{{\eta}})$ unchanged. To compute the expected value of the log-likelihood term, we use Monte-Carlo integration by drawing samples from the variational distribution $q$ given the parameters $\bm{{\eta}}$. The likelihood function is computed using Section˜3.2 and we approximate the first term in Eq.˜15, as follows:
@@
$=\frac{1}{N_{q}}\sum_{c=1}^{N_{q}}\log\left(\prod_{i=1}^{N_{1}}\prod_{k=1}^{N_{2}}\mathcal{N}\left(s_{k}^{(i)};\mu_{k}\left(\bm{{\theta}}^{(c)}\right),\sigma_{k}\left(\bm{{\theta}}^{(c)}\right)\right)\right),\hskip 11.38092pt\bm{{\theta}}^{(c)}\sim q(\bm{{\theta}};\bm{\eta})$
$=\frac{1}{N_{q}}\sum_{c=1}^{N_{q}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\log\left(\mathcal{N}\left(s^{(i)}_{k};\mu_{k}\left(\bm{{\theta}}^{(c)}\right)\sigma_{k}\left(\bm{{\theta}}^{(c)}\right)\right)\right),\hskip 11.38092pt\bm{{\theta}}^{(c)}\sim q(\bm{{\theta}};\bm{\eta}).$ (16)
where $N_{q}$ is the number of Monte Carlo samples, and $\bm{{\theta}}^{(c)}$ is the $c^{\text{th}}$ sample drawn from $q(\bm{{\theta}};\bm{\eta})$. In practice, we found that using 25 Monte Carlo samples ($N_{q}=25$) is sufficient for accurate approximation, aligning with findings in [kucukelbir2017automatic] which suggest that even a single sample can yield satisfactory results. This follows the common practice of using a fixed number of samples in variational inference [rezende2014stochastic, kingma2013auto]. Alternatively, adaptive sampling or variance reduction techniques have also been proposed to mitigate gradient variance, such as control variates and adaptive reparameterisation methods [roeder2017sticking]. In the next section, we detail Rényi’s $\alpha$-divergence as an alternative to KLD, which represents the second term of our objective function.
3.4 Rényi’s $\alpha$-divergence
Rényi’s $\alpha$-divergence, introduced by Rényi [renyi1961measures], offers a robust alternative to KLD for addressing prior misspecification. It is characterised by a hyperparameter $\alpha$, which controls the degree of robustness, and it converges to KLD as a limiting case when $\alpha$ approaches 1. We denote this divergence by $D_{AR}^{(\alpha)}$ and follow the parameterisation provided in [cichocki2010families]:
$D_{AR}^{(\alpha)}[q(\theta)\mid\mid p(\theta)]=\frac{1}{\alpha(\alpha-1)}\log\left(\mathbb{E}_{q(\theta)}\left[\left(\frac{p(\theta)}{q(\theta)}\right)^{1-\alpha}\right]\right).$ (17)
@@ 3.4 Rényi’s $\alpha$-divergence
In the context of our problem, $D_{AR}^{(\alpha)}$ represents the second term of the objective function in Eq.˜15, which is approximated as follows:
$D_{AR}^{(\alpha)}[q(\bm{{\theta}};\bm{\eta})\mid\mid p(\bm{{\theta}})]\approx\frac{1}{\alpha(\alpha-1)}\log\left(\frac{1}{N_{q}}\sum_{c=1}^{N_{q}}\left[\left(\frac{p(\bm{{\theta}}^{(c)}))}{q(\bm{{\theta}}^{(c)};\bm{\eta})}\right)^{1-\alpha}\right]\right),$ (19)
@@
3.4.1 Mathematical mechanism of prior robustness
The robustness of Rényi’s $\alpha$-divergence to prior misspecification arises from how it balances the influence of prior beliefs versus observed data, with the behaviour depending critically on the value of $\alpha$.
$\bm{\alpha\in(0,1)}$ – reduces prior constraint: The divergence contains the term $(p(\bm{{\theta}})/q(\bm{{\theta}}))^{(1-\alpha)}$, which creates an asymmetric penalty structure that promotes robustness to prior misspecification. When the posterior $q(\bm{{\theta}})$ has mass in regions where the prior $p(\bm{{\theta}})$ has little support ($p\approx 0$), the ratio $p/q\approx 0$, making $(p(\bm{{\theta}})/q(\bm{{\theta}}))^{(1-\alpha)}\approx 0$ and contributing minimal penalty. This allows the posterior to explore data-supported regions even when they contradict the prior. Conversely, when the posterior has little mass in prior-supported regions ($q\approx 0$ but $p>0$), the ratio p/q becomes large, creating a substantial penalty that prevents the posterior from completely ignoring the prior. This asymmetric behaviour naturally leads to wider posterior distributions (larger marginal variances) that provide more conservative uncertainty estimates and prevent overconfidence in parameter values influenced by incorrect prior assumptions.
$\bm{{\alpha>1}}$ – increases data focus: When $\alpha>1$, the divergence provides the opposite behaviour: it creates more concentrated posterior distributions (smaller marginal variances) that focus more heavily on the data-fit term. This effectively reduces the influence of the prior more aggressively, moving the posterior closer to what would be obtained by maximum likelihood estimation alone. This can be beneficial when the prior is known to be poorly specified and the goal is to rely primarily on the data. This dual capability makes Rényi’s $\alpha$-divergence particularly valuable for BNNs, where specifying meaningful priors for thousands of parameters is challenging. Standard approaches often use simple factorised priors (e.g., independent Gaussians) that may not reflect the true parameter relationships. The $\alpha$-divergence allows practitioners to either maintain broader uncertainty ($\alpha$ < 1) or focus more aggressively on data-driven solutions ($\alpha$ > 1), depending on their confidence in the prior specification.
@@ 3.4.1 Mathematical mechanism of prior robustness
In Fig.˜3, we plot divergence values $D_{AR}^{(\alpha)}$ between $q(\theta)$ and $p(\theta)$, assuming both are normal distributions with $p(\theta)$ as a standard normal distribution. The divergence value decreases as $\alpha$ increases, with KLD as the case when $\alpha=1$.
@@
The behaviour of the optimisation problem differs significantly depending on whether $\alpha$ is greater than or less than 1, with important implications for robustness to prior misspecification. For $\alpha>1$, $D_{AR}^{(\alpha)}$ exhibits a mode-seeking behaviour, causing the posteriors to become more concentrated while being penalised less for deviating from the prior distribution. In this regime, the objective function places more emphasis on the data-loss term, moving the posterior closer to the empirical risk minimiser and reducing the influence of potentially misspecified priors. Conversely, for $\alpha\in(0,1)$, $D_{AR}^{(\alpha)}$ encourages a mass-covering behaviour and provides robustness to contradictory priors by penalising deviations from the prior less heavily than KLD. This allows the posterior to maintain larger marginal variances and spread its mass in a less concentrated way, preventing it from being unduly influenced by a flawed prior. As $\alpha$ decreases below 1, the values of $D_{AR}^{(\alpha)}$ increase, but importantly, this larger regularisation term does not necessarily imply that a misspecified prior will dominate the inference, as the structural properties of Rényi’s $\alpha$-divergence provide inherent robustness to prior misspecification [knoblauch2019generalized].
3.4.2 Choice of $\alpha$
@@
Given that the true posterior is unknown, it is uncertain whether the mode-seeking or mass-covering behaviour is preferable, making the choice problem-dependent. For scenarios where the true posterior is multimodal, a divergence that produces wider parameter variances (i.e., mass-covering behaviour) may be beneficial for achieving more robust predictive estimates. In particular, we aim to set $D_{AR}^{(\alpha)}$ such that the posterior is robust to priors that strongly contradict observed data, and provides reliable uncertainty quantification. Similar to [li2016renyi, knoblauch2019generalized], we rely on validation sets and adjust $\alpha$ as a hyperparameter to minimise metrics such as the normalised mean squared error or negative log-likelihood.
In this work, we adopt a grid-based search over a specified range of $\alpha$ values, evaluating model performance on held-out cross-validation data. This allows us to empirically assess the trade-off between predictive accuracy and uncertainty calibration induced by different settings of $\alpha$. Based on our experiments across diverse problems, an initial range of $\alpha$ between 0.50 and 2.00 serves as a reasonable starting point, as it encompasses the values that consistently yielded competitive performance. Smaller values ($\alpha$ < 1) may be beneficial for problems where uncertainty quantification demands broader mass coverage, whereas larger values ($\alpha$ > 1) may be preferable for problems where precise mode-seeking behaviour is needed. These observations align with previous findings in [knoblauch2019generalized].
While this grid-based strategy is less efficient than automated hyperparameter optimisation, it was chosen deliberately to provide a systematic mapping of the $\alpha$-performance landscape. This mapping reveals how different regimes ($\alpha<1$ versus $\alpha>1$) influence predictive accuracy and uncertainty calibration, offering interpretability that direct optimisation alone would obscure. In practice, however, our framework is fully compatible with alternative selection strategies. Standard hyperparameter optimisation techniques, such as Bayesian optimisation (BO), are straightforward to integrate and are expected to identify competitive $\alpha$ values with far fewer evaluations. Moreover, gradient-based or meta-learning approaches hold promise for dynamically adapting $\alpha$ during training, an exciting direction for improving both efficiency and adaptability in future work.